one canned juice drink is 15% orange juice, another is 5% orange juice. How many liters of each should be mixed together in order to get 10L that is 14% orange juice?
Answer by Alan3354(69443)

one canned juice drink is 15% orange juice, another is 5% orange juice. How many liters of each should be mixed together in order to get 10L that is 14% orange juice?
--------------------
f = 15%
v = 5%
0.15f + 0.05v = 0.14*10
15f + 5v = 140
3f + v = 28
f + v = 10 (total volume)
---------- Subtract
2f = 18
f = 9 liters of 15%
v = 1 liter of 5%